OpenStudy (kendricklamar2014):

What is the value of the range of the function f(x)=x^2+2 for the domain value 1/4? A. Y = 2 B. Y = 2 1/4 C. Y = 2 1/8 D. Y = 2 1/16

OpenStudy (sidsiddhartha):

domain means all the x-values, and the range is all the y-values. so domain =x=1/4 now just plug x=1/4 in the equation to find f(x) and that will be your range

OpenStudy (sidsiddhartha):

which will produce \[Y=f(1/4)=(\frac{ 1 }{ 4 })^2+2=1/16+2=33/16=2\frac{ 1 }{ 16 }\]

OpenStudy (kendricklamar2014):

so what do u do next @sidsiddhartha

OpenStudy (sidsiddhartha):

ive already solved it D. Y = 2 1/16 thATS the answer ,ok?

OpenStudy (kendricklamar2014):

thanks

OpenStudy (sidsiddhartha):

if u have a x value then just put it in the equation in place of x to get y value

OpenStudy (sidsiddhartha):

yw!!